OpenStudy (anonymous):

PLEASE HELP!!!! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

You are given the equation 2x+3y=1200, so all you want to do is solve for y in this equation to get it into the form y=mx+b. So you would do... 2x+3y=1200 3y=-2x+1200 (3/3)y= (-2/3)x+(1200/3) y=(-2/3)x+400 Now you can probably recognize that y=(-2/3)x+400 is in y=mx+b form. So the slope of y=mx+b is the m value, and the y-intercept is the b value. Therefore the slope of this equation is -2/3 and the y-intercept is 400. I hope this helps!

OpenStudy (anonymous):

THANK YOU!!!! :)

OpenStudy (anonymous):

yw